What times 4/5 equels 2/5
Answer by candydreams99(2)

Answer : ½

4/5 x N = 2/5
Multiply throughout by reciprocal of 1/5(which should look like)
5(4/5) x N = 2/5 (5)
4N = 2
N = 2/4
N = 1/2